Given the following information about copper:

 

Specific heat of solid 0.385 J/(g⋅°C)

Specific heat of liquid 0.490 J/(g⋅°C)

Melting point 1086°C

Boiling point 2570°C

Heat of fusion 205 J/g

Heat of vaporization 4799 J/g 


Which quantity of energy is involved in converting 1.50 g of solid copper at 25.0°C to a liquid at 2000.°C?

1
Expert's answer
2021-04-07T11:28:00-0400

E1 = mc∆T

T1 = 298 K , T2 = 373 K , "\Delta T" = 373-298 = 75K

E1 = 1.50×4.18×75 = 470.25 joules

heat of vaporization E2 = 1.5 ×4799 joulees = 7198.5 joules

E3 = mc∆T

T2 = 373 K, T3 = 2273K ,∆T = 2273-373 =1900K

E3 = 1.5×4.18×1900 =11,913joules

Total Energy (E) = E1 + E2 + E3

E = 470.25 + 7198.5 + 11,913

E = 19,581.25 joules